# Gibbs polystability of Fano manifolds, stability thresholds and symmetry breaking

### v1

@@ 3.3.1. Comparison with Tian’s orthogonality condition
for all $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ satisfying the orthogonality condition 3.29.222Strictly speaking, in the original conjecture [89, Conj 5.5] the role of $\mathcal{M}$ is played by the Ding functional $\mathcal{D}$ (see Section 4.0.1); anyhow, both versions of the conjecture were established in [48]. It should be emphasized that the functional $(\mathcal{I}-\mathcal{J})$ appearing in the right hand side above may as well be replaced by the functionals $\mathcal{I}$ or $\mathcal{J}$ (up to changing the constant $\epsilon),$ since all three functionals $\mathcal{I},(\mathcal{I}-\mathcal{J})$ and $\mathcal{J}$ are comparable (up to multiplicative constants [23, Section 2]). The inequality 3.30 should be compared to item 3 in Theorem 3.7, saying that the inequality 3.30 holds for all $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ such that $\omega_{u}^{n}\boldsymbol{m}$ integrates to zero on $X.$ Note that, while Tian’s orthogonality condition 3.29 is linear in $u$, the moment condition appearing in Theorem 3.7 is linear with respect to the Monge–Ampère measure $\omega_{u}^{n}.$ The relation between these to conditions is clarified in the following lemma, which gives a refinement of [95, Lemma A.2]:
@@
where $\Delta$ denotes the Laplacian with respect to the Kähler metric $\omega_{0}.$ Moreover, there exists a unique $v\in u\cdot G\subset\mathcal{H}(X,\omega_{0})/\mathbb{R}$ satisfying any of these two conditions. When $\omega_{0}$ is Kähler–Einstein the condition $(ii)$ is equivalent to Tian’s orthogonality condition 3.29 for $v.$
@@
using, in the first equality, that the Futaki character vanishes (formula 3.15). Expressing $\left\langle dv,JW\right\rangle$ as the scalar product $\nabla v\cdot JW$ between the vector field $JW$ and the gradient $\nabla v$ of $v$ with respect to the Kähler metric $\omega_{0}$. Using that $JW$ is proportional to $\nabla h_{W},$ where $h_{W}$ is the Hamiltonian function corresponding to $W,$ we deduce that
@@
Hence, $\frac{d}{dt}\mathcal{J}(v_{t})=0$ for all $W\in\mathfrak{k}$ iff condition $(ii)$ holds, as desired. Finally, when $\omega_{0}$ is Kähler–Einstein $\Delta h_{W}=\lambda_{1}h_{W}$ for any $W\in\mathfrak{k},$ where $\lambda_{1}$ denotes the smallest strictly positive eigenvalue of $\Delta$ [8, Thm 2.2]. Since $\boldsymbol{m}=(h_{W_{1}},\dots,h_{W_{r}})$ for a basis $W_{i}$ in $\mathfrak{k}$ it follows that $(ii)$ is equivalent to Tian’s condition 3.29, when $\omega_{0}$ is Kähler–Einstein. ∎
In the case when $n=1,$ so that $X=\mathbb{P}^{1}$, the functionals $\mathcal{I}-\mathcal{J}$ and $\mathcal{J}$ coincide and thus the moment condition appearing in Theorem 3.7 then coincides with condition $(ii)$ in the previous lemma. The formula $\delta^{A}(\mathbb{P}^{1},0)^{\mathcal{G}}=2$ that we shall establish in the next section thus yields a quasi-sharp inequality of the form 3.30 on $\mathbb{P}^{1},$ saying that the inequality holds for any $\epsilon\in]0,1[$ - but not for $\epsilon>1$ - when $u$ satisfies Tian’s orthogonality condition 3.29.
@@ 3.4. The reduced analytic stability thresholds of K-polystable log Fano curves
Consider the case when the fixed $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)})$ is induced by the Fubini–Study metric on $\mathcal{O}(1)\rightarrow\mathbb{P}^{1}$ and thus $\mathrm{SU}(2)$–invariant. Then, identifying $\mathbb{P}^{1}$ with the unit-sphere in $\mathbb{R}^{3},$ $\omega_{0}=V\omega$ for the unique $\mathrm{SO}(3)$–invariant probability measure $\omega$ and
$E(\mu)=-\frac{V}{2}\int_{X\times X}\log\left(|x-y|^{2}\right)\mu(x)\otimes\mu(y)+C$
for some constant $C$ (by Example 3.2, using the explicit Green function for $(X,\omega_{0})$ ). Moreover, when $\Delta=0$ the measure $\mu_{0}$ is proportional to $\omega_{0}$ and $\boldsymbol{m}=\boldsymbol{x}.$ When $w\neq 0,$ taking $x_{0}$ and $x_{1}$ as the south and north pole on the unit sphere, $\mu_{0}$ is proportional to $e^{-|\boldsymbol{x}-\boldsymbol{x}_{0}|^{2}-|\boldsymbol{x}-\boldsymbol{x}_{1}|^{2}}\omega_{0}$ and $\boldsymbol{m}=2x_{3}/V.$
@@
where $\phi_{\mu}$ is the potential of $\mu$ (formula 3.7).
@@
for $r$ sufficiently small (as follows, for example, from Green’s formula in $\mathbb{C}$ or Skoda’s integrability theorem in $\mathbb{C}^{n}$). Now, first consider the case $w=0$ and take $\mu\in\mathcal{P}(X)_{0}.$ By the previous lemma, $\mu(\left\{x\right\})\leq 1/2$ for any $x\in X.$ Since $V=2$ this means that $dd^{c}\phi_{\mu}(\left\{x\right\})\leq 1$ for any $x\in X.$ By 3.32, this implies that $e^{-t\phi_{\mu}}\in L_{\mathrm{loc}}^{1}$ for any $t<1.$ Hence, $\alpha(\mathbb{P}^{1},0)_{0}\geq 1.$ Finally, in the case $w\neq 0$ we may assume that $\Delta_{w}=w([0]+[\infty]).$ On a disc $D_{r}$ of radius $r$ centered at $x=0,$
@@
Hence, by 3.32, $\int_{D_{r}(0)}e^{-t(\phi-\phi_{0})}\mu_{0}<\infty$ if $t<1$ and the same argument applies when 0 is replaced by $\infty.$ Finally, for $x$ different from 0 and $\infty$ we can use that $(dd^{c}\psi)(\{x\})\leq tV$ (since $\int dd^{c}\phi=V$) to deduce that $\int_{D_{r}(x)}e^{-t(\phi-\phi_{0})}\mu_{0}<\infty$ when t<1/V=1/(22w),t<1/V=1/(2-2w), as desired. ∎
@@
where $F(-\gamma)_{0}$ is defined by formula 3.13. Moreover, $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ and $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}=\min\left\{\frac{1}{1-w},2\right\}$ when $w\neq 0.$
@@
Let us first show that for any log Fano curve $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}\geq 2\alpha(X,\Delta)_{0}.$ We will adapt the argument in [10] (concerning the case $\mathcal{G}$ trivial) to the present setup. Fix $\gamma>0$ and take $\mu\in\mathcal{P}(X)_{0}.$ Using that $\gamma u_{\mu}$ is a candidate for the sup appearing in the Legendre duality formula 3.14 for $D(\mu)$ gives, for any $t<\alpha(X,\Delta)_{0}$
@@
using Lemma 3.12. Next, setting $u_{\mu}\coloneqq\phi_{\mu}-\phi_{0}$ (formula 3.7) we deduce, using $E(\mu)=-\int_{X}u_{\mu}\mu/V2$ (Example 3.2),
@@
where $v^{\prime\prime}(y)$ denotes the absolutely continuous part of the second distributional derivatives $\partial^{2}v$ and $v^{\prime}$ is the function in $L_{\text{loc}}^{\infty}$ defined by the first distributional derivative $\partial v$ (Prop 8.1, combined with 8.3). In particular,
@@
which tends to -$\infty$ if $\gamma>2,$ proving the inequality 3.34. It remains to show that when $w\in]0,1[$
@@
The corresponding Legendre transform $v_{t}$ is given by $v_{t}(y)=v_{0}(y)-ty.$ Hence, by formula 8.2 in the appendix, we get as before (but now using that $\mathcal{L}(v_{t})=0)$
@@ 3.5. Stability thresholds on the space of geodesic rays vs divisorial valuations
For the log Fano curves $(X,\Delta_{w})$ we have $\overrightarrow{\delta}_{0}(X,0)=2$ and $\overrightarrow{\delta}_{0}(X,\Delta_{w})=\min\{\frac{1}{1-w},2\}$ when $w\neq 0.$ Moreover, when $\phi_{0}$ is induced by the Fubini–Study metric, there is a geodesic ray $u_{t}$ in $\boldsymbol{H}_{1,1}(X,\omega_{0})_{0},$ attaining the inf defining $\overrightarrow{\delta}_{0}(X,\Delta)$, with the property that the limit of $\omega_{u_{t}}/V$ in $\mathcal{P}(X)_{0}$ as $t\rightarrow\infty$ is the measure $\mu$ defined by formula 3.33.
@@
The geodesic ray $u_{t}$ appearing in the previous proposition is attached to a test configuration $(\mathcal{X},\mathcal{L})$ for $(X,-K_{(X,\Delta)})$ and thus defines a psh metric on the Berkovich analytification of $(X,-K_{(X,\Delta)}),$ with respect to the trivial valuation on $\mathbb{C},$ in the sense of [33, 35]. Moreover, the Monge–Ampère measure $\text{MA(}(\mathcal{X},\mathcal{L}))$ of $(\mathcal{X},\mathcal{L})$ on the Berkovich analytification $X^{\text{an}}$ of $X$ coincides with the measure $\mu$ defined by formula 3.33, under the standard embedding of a complex curve $X$ into $X^{\text{an}}$ (where $x\in X$ is mapped to the divisorial valuation defined by $x).$ This follows from the fact that, by construction, the test configuration $(\mathcal{X},\mathcal{L})$ is toric for some $\mathbb{C}^{*}$–action on $\mathbb{P}^{1}$. Thus formula 3.38 translates into the following weak convergence of measures on the Berkovich analytification of $\mathbb{P}^{1}$:
@@
4. Applications to logarithmic Hardy–Littlewood–Sobolev inequalities, Moser–Trudinger inequalities and spontaneous symmetry breaking
@@
$E(\mu)\leq\frac{1}{\gamma}D(\mu|\mu_{0})+C_{\gamma}$
for any $\mu\in\mathcal{P}(X).$ When $w=0$ and $\gamma=1$ this inequality is known as the logarithmic Hardy–Littlewood–Sobolev (HLS) inequality on the two-sphere $X$ [37, 38]. In this case it holds without the moment constraint (since $F_{-1}(\mu)$ is $\mathcal{G}$–invariant). Thus Theorem 3.13 yields an improvement of the logarithmic HLS inequality, allowing any $\gamma\in]0,2[,$ when $\mu$ has vanishing moment (formula 3.1). In contrast to [37, 38], it does not seem possible to use symmetric decreasing rearrangements when $\gamma>1$ (see Section 4.1.1).
We next show that Theorem 3.13 also implies Aubin’s improvement of the Moser–Trudinger inequality on $S^{2}$ [7, Cor 2]. To this end, first recall that, given a log Fano manifold $(X,\Delta)$ and a reference metric $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)}),$ the twisted Ding functional on $\mathcal{H}(X,\omega_{0})$ is defined by
@@
It was introduced in [51] (when $\Delta=0)$ [^3] In the notation of [51], $\mathcal{D}_{\gamma}(u)=F_{\gamma}(u),$ but here the notation $F_{\gamma}$ is reserved for the free energy functional.. Here we will consider the case of a log Fano curve and allow $u\in C^{1}(X)$ (in particular, we do not assume $\omega_{u}\geq 0).$ In this case $\mathcal{E}(u)$ is given by the classical expression in Example 3.2. We will restrict $\mathcal{D}_{\gamma}(u)$ to the subspace $C^{1}(X)_{0}$ of all $u\in C^{1}(X)$ satisfying
@@
This is essentially shown in [10], but for completeness we recall the proof. The starting point is that when $n=1$, the sup in the definition 3.5 of $E(\mu)$ may be taken over all of $C^{0}(X),$ as pointed out in Example 3.2. Hence, $E(\mu)\leq\left\langle v,\mu\right\rangle-\mathcal{E}(v)$ for any $v\in C^{0}(X).$ Now take $\mu=e^{-\gamma u}/\int_{X}e^{-\gamma u}\mu_{0}$ and apply the previous inequality for $v=u.$ The inequality 4.3 then follows from the identity
@@
which is a direct consequence of the definition of $D(\mu|\mu_{0}).$ Thanks to the inequality 4.3 the lower bound on $\mathcal{D}_{\gamma}$ in the corollary thus follows directly from the lower bound on $\delta^{A}(X,\Delta_{w})^{\mathcal{G}}$ in Theorem 3.13. Next, to destabilize $\mathcal{D}_{\gamma}(u),$ first recall [10] that for any $u\in\mathcal{H}(X,\omega_{0})$
@@
In fact, this inequality holds also when $n>1$ by replacing $\omega_{u}$ with $\omega_{u}^{n}.$ Indeed, $F_{\gamma}(\frac{\omega_{u}}{V})-\mathcal{D}_{\gamma}(u)=D(\frac{\omega_{u}^{n}}{V}|\frac{\mu_{\phi_{0}+u}}{\int\mu_{\phi_{0}+u}})\geq 0,$ using that $D(\cdot|\cdot)\geq$ on $\mathcal{P}(X)^{2}.$ All that remains is thus to show that the curve $\mu_{t}$ in $\mathcal{P}(X)_{0}$ destabilizing $F_{-\gamma},$ constructed in the proof of Theorem 3.13, has the property that it can be expressed as $\mu_{t}=\omega_{u_{t}}/V$ for $u_{t}$ in $\in C^{1,1}(X)_{0}\cap\text{PSH}(X,\omega_{0}).$ To this end first consider the case when $w<1/2,$ i.e. $\min\{\frac{1}{1-w},2\}<2.$ Then $u_{t}:=\phi_{t}-\phi_{0},$ where $\phi_{t}$ is defined by 3.37, has the desired properties. Indeed, $\mu_{t}$ destabilizes $F_{-\gamma}$ on $\mathcal{P}(X)_{0}$ and, just as before, the $\mathbb{Z}_{2}$–invariance of $u_{t}$ (resulting from the $\mathbb{Z}_{2}$–invariance of $\phi_{t}$ and $\phi_{0})$ implies that $e^{-\gamma u_{t}}\mu_{0}$ has vanishing moment. Finally, when $w\geq 1/2$ the curve $u_{t}:=\phi_{t}-\phi_{0},$ where $\phi_{t}$ is the Legendre transform of $v_{t},$ defined by formula 3.35, has the desired properties. Again, this follows from the symmetry argument employed in the proof of Theorem 3.13, since $u_{t}$ is $\mathbb{Z}_{2}$–invariant. ∎
@@
is equivalent to the following Moser–Trudinger inequality with Aubin’s constraint:

### v2

@@ 3.3.1. Comparison with Tian’s orthogonality condition
for all $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ satisfying the orthogonality condition 3.31.222Strictly speaking, in the original conjecture [89, Conj 5.5] the role of $\mathcal{M}$ is played by the Ding functional $\mathcal{D}$ (see Section 4.0.1); anyhow, both versions of the conjecture were established in [48]. It should be emphasized that the functional $(\mathcal{I}-\mathcal{J})$ appearing in the right hand side above may as well be replaced by the functionals $\mathcal{I}$ or $\mathcal{J}$ (up to changing the constant $\epsilon),$ since all three functionals $\mathcal{I},(\mathcal{I}-\mathcal{J})$ and $\mathcal{J}$ are comparable (up to multiplicative constants [23, Section 2]). The inequality 3.32 should be compared to item 3 in Theorem 3.7, saying that the inequality 3.32 holds for all $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ such that $\omega_{u}^{n}\boldsymbol{m}$ integrates to zero on $X.$ Note that, while Tian’s orthogonality condition 3.31 is linear in $u$, the moment condition appearing in Theorem 3.7 is linear with respect to the Monge–Ampère measure $\omega_{u}^{n}.$ The relation between these to conditions is clarified in the following lemma, which gives a refinement of [95, Lemma A.2]:
@@
where $\Delta$ denotes the Laplacian with respect to the Kähler metric $\omega_{0}.$ Moreover, there exists a unique $v\in u\cdot G\subset\mathcal{H}(X,\omega_{0})/\mathbb{R}$ satisfying any of these two conditions. When $\omega_{0}$ is Kähler–Einstein the condition $(ii)$ is equivalent to Tian’s orthogonality condition 3.31 for $v.$
@@
using, in the first equality, that the Futaki character vanishes (formula 3.17). Expressing $\left\langle dv,JW\right\rangle$ as the scalar product $\nabla v\cdot JW$ between the vector field $JW$ and the gradient $\nabla v$ of $v$ with respect to the Kähler metric $\omega_{0}$. Using that $JW$ is proportional to $\nabla h_{W},$ where $h_{W}$ is the Hamiltonian function corresponding to $W,$ we deduce that
@@
Hence, $\frac{d}{dt}\mathcal{J}(v_{t})=0$ for all $W\in\mathfrak{k}$ iff condition $(ii)$ holds, as desired. Finally, when $\omega_{0}$ is Kähler–Einstein $\Delta h_{W}=\lambda_{1}h_{W}$ for any $W\in\mathfrak{k},$ where $\lambda_{1}$ denotes the smallest strictly positive eigenvalue of $\Delta$ [8, Thm 2.2]. Since $\boldsymbol{m}=(h_{W_{1}},\dots,h_{W_{r}})$ for a basis $W_{i}$ in $\mathfrak{k}$ it follows that $(ii)$ is equivalent to Tian’s condition 3.31, when $\omega_{0}$ is Kähler–Einstein. ∎
In the case when $n=1,$ so that $X=\mathbb{P}^{1}$, the functionals $\mathcal{I}-\mathcal{J}$ and $\mathcal{J}$ coincide and thus the moment condition appearing in Theorem 3.7 then coincides with condition $(ii)$ in the previous lemma.
@@ 3.4. The reduced analytic stability thresholds of K-polystable log Fano curves
Consider the case when the fixed $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)})$ is induced by the Fubini–Study metric on $\mathcal{O}(1)\rightarrow\mathbb{P}^{1}$ and thus $\mathrm{SU}(2)$–invariant. Then, identifying $\mathbb{P}^{1}$ with the unit-sphere in $\mathbb{R}^{3},$ $\omega_{0}=V\omega$ for the unique $\mathrm{SO}(3)$–invariant probability measure $\omega$ on $S^{2}.$ Moreover, when $\Delta=0$ the measure $\mu_{0}$ is proportional to $\omega_{0}$ and $\boldsymbol{m}=\boldsymbol{x}.$ When $w\neq 0,$ taking $x_{0}$ and $x_{1}$ as the south and north pole on the unit sphere, $\mu_{0}$ is proportional to $|\boldsymbol{x}-\boldsymbol{x}_{0}|^{-2w}|\boldsymbol{x}-\boldsymbol{x}_{1}|^{-2w}\omega_{0}$ and $\boldsymbol{m}=2x_{3}/V.$
@@
where $\phi_{\mu}$ is the potential of $\mu$ (formula 3.8).
@@
for $r$ sufficiently small (as follows, for example, from Green’s formula in $\mathbb{C}$ or Skoda’s integrability theorem in $\mathbb{C}^{n}$). Now, first consider the case $w=0$ and take $\mu\in\mathcal{P}(X)_{0}.$ By the previous lemma, $\mu(\left\{x\right\})\leq 1/2$ for any $x\in X.$ Since $V=2$ this means that $dd^{c}\phi_{\mu}(\left\{x\right\})\leq 1$ for any $x\in X.$ By 3.34, this implies that $e^{-t\phi_{\mu}}\in L_{\mathrm{loc}}^{1}$ for any $t<1.$ Hence, $\alpha(\mathbb{P}^{1},0)_{0}\geq 1.$ Finally, in the case $w\neq 0$ we may assume that $\Delta_{w}=w([0]+[\infty]).$ On a disc $D_{r}$ of radius $r$ centered at $x=0,$
@@
Hence, by 3.34, $\int_{D_{r}(0)}e^{-t(\phi-\phi_{0})}\mu_{0}<\infty$ if $t<1$ and the same argument applies when 0 is replaced by $\infty.$ Finally, for $x$ different from 0 and $\infty$ we can use that $(dd^{c}\psi)(\{x\})\leq tV$ (since $\int dd^{c}\phi=V$) to deduce that $\int_{D_{r}(x)}e^{-t(\phi-\phi_{0})}\mu_{0}<\infty$ when t<1/V=1/(22w),t<1/V=1/(2-2w), as desired. ∎
@@
where $F(-\gamma)_{0}$ is defined by formula 3.15. Moreover, $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ and $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}=\min\left\{\frac{1}{1-w},2\right\}$ when $w\neq 0.$
@@
Let us first show that for any log Fano curve $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}\geq 2\alpha(X,\Delta)_{0}.$ We will adapt the argument in [10] (concerning the case $\mathcal{G}$ trivial) to the present setup. Fix $\gamma>0$ and take $\mu\in\mathcal{P}(X)_{0}.$ Using that $\gamma u_{\mu}$ is a candidate for the sup appearing in the Legendre duality formula 3.16 for $D(\mu)$ gives, for any $t<\alpha(X,\Delta)_{0}$
@@
using Lemma 3.12. Next, setting $u_{\mu}\coloneqq\phi_{\mu}-\phi_{0}$ (formula 3.8) we deduce, using $E(\mu)=-\int_{X}u_{\mu}\mu/V2$ (Example 3.2),
@@
where $v^{\prime\prime}(y)$ denotes the absolutely continuous part of the second distributional derivatives $\partial^{2}v$ and $v^{\prime}$ is the function in $L_{\text{loc}}^{\infty}$ defined by the first distributional derivative $\partial v$ (Prop 9.1, combined with 9.3). In particular,
@@
which tends to -$\infty$ if $\gamma>2,$ proving the inequality 3.36. It remains to show that when $w\in]0,1[$
@@
The corresponding Legendre transform $v_{t}$ is given by $v_{t}(y)=v_{0}(y)-ty.$ Hence, by formula 9.2 in the appendix, we get as before (but now using that $\mathcal{L}(v_{t})=0)$
@@ 3.5. Stability thresholds on the space of geodesic rays vs divisorial valuations
For the log Fano curves $(X,\Delta_{w})$ we have $\overrightarrow{\delta}_{0}(X,0)=2$ and $\overrightarrow{\delta}_{0}(X,\Delta_{w})=\min\{\frac{1}{1-w},2\}$ when $w\neq 0.$ Moreover, when $\phi_{0}$ is induced by the Fubini–Study metric, there is a geodesic ray $u_{t}$ in $\boldsymbol{H}_{1,1}(X,\omega_{0})_{0},$ attaining the inf defining $\overrightarrow{\delta}_{0}(X,\Delta)$, with the property that the limit of $\omega_{u_{t}}/V$ in $\mathcal{P}(X)_{0}$ as $t\rightarrow\infty$ is the measure $\mu$ defined by formula 3.35.
@@
The geodesic ray $u_{t}$ appearing in the previous proposition is attached to a test configuration $(\mathcal{X},\mathcal{L})$ for $(X,-K_{(X,\Delta)})$ and thus defines a psh metric on the Berkovich analytification of $(X,-K_{(X,\Delta)}),$ with respect to the trivial valuation on $\mathbb{C},$ in the sense of [33, 35]. Moreover, the Monge–Ampère measure $\text{MA(}(\mathcal{X},\mathcal{L}))$ of $(\mathcal{X},\mathcal{L})$ on the Berkovich analytification $X^{\text{an}}$ of $X$ coincides with the measure $\mu$ defined by formula 3.35, under the standard embedding of a complex curve $X$ into $X^{\text{an}}$ (where $x\in X$ is mapped to the divisorial valuation defined by $x).$ This follows from the fact that, by construction, the test configuration $(\mathcal{X},\mathcal{L})$ is toric for some $\mathbb{C}^{*}$–action on $\mathbb{P}^{1}$. Thus formula 3.40 translates into the following weak convergence of measures on the Berkovich analytification of $\mathbb{P}^{1}$:
@@
4. Applications to logarithmic HLS inequalities, Moser–Trudinger inequalities and spontaneous symmetry breaking
@@
(4.1) $E(\mu)\leq\frac{1}{\gamma}D(\mu|\mu_{0})+C_{\gamma}\,\,\forall\mu\in\mathcal{P}(X):\int_{X}\boldsymbol{m}\mu=0.$
When $w=0$ and $\gamma=1$ this inequality is known as the logarithmic Hardy–Littlewood–Sobolev (HLS) inequality on the two-sphere $X$ [37, 38, 9]. In this case it holds without the moment constraint (since $F_{-1}(\mu)$ is $\mathcal{G}$–invariant). Thus Theorem 3.13 yields an improvement of the logarithmic HLS inequality, allowing any $\gamma\in]0,2[,$ under the constraint that $\mu$ has vanishing moment. In contrast to [37, 38, 9], in this case it does not seem possible to use symmetric decreasing rearrangements, due to a phenomenon of spontaneous symmetry breaking (see Section 4.1.1).
To facilitate the comparison between our complex-geometric notation and the PDE-literature [37, 39, 38, 9, 72] we provide the following basic lemma, proved in the last section of the appendix:
Lemma 4.1.
(normalizations) Let $\omega_{\text{$FS$}}$ denote the Fubini-Study Kähler form on $\mathbb{P}^{1}$ (which is $SU(2)$-invariant and thus coincides with the $SO(3)$-invariant probability measure on $S^{2})$ and denote by $\Delta_{\text{FS}}$ the Laplacian on $\mathbb{P}^{1}$ defined by
(4.2) $dd^{c}u=\Delta_{\text{FS}}u\omega_{\text{FS}},$
Then $\Delta_{\text{FS}}$ coincides with the Laplacian $\Delta_{g_{S^{2}}}$ wrt the metric $g_{S^{2}}$ on $S^{2}$ induced by the Euclidean metric on $\mathbb{R}^{3}.$ Moreover, when $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta_{w})})$ is $SU(2)$-invariant, i.e. when $\omega_{0}=V\omega_{\text{FS}}$ for V=(22w),V=(2-2w),
(4.3) $E(\mu)=(1-w)\int_{S^{2}\times S^{2}}G\mu\otimes\mu,$
where $G(\boldsymbol{x},\boldsymbol{y})$ is the Green function for $\Delta_{g_{S^{2}}},$ i.e. $G(\boldsymbol{x},\boldsymbol{y})=\log\left(\frac{|\boldsymbol{x}-\boldsymbol{y}|^{2}}{4}\right)+1.$
In Section 5 we will establish a sharp version of the constrained logarithmic HLS inequality 4.1 when $w=0.$ But first, in this section we show that Theorem 3.13 implies a generalization of Aubin’s improvement of the Moser–Trudinger inequality on $S^{2}$ [7, Cor 2] to $w\neq 0.$ To this end, first recall that, given a log Fano manifold $(X,\Delta)$ and a reference metric $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)}),$ the twisted Ding functional on $\mathcal{H}(X,\omega_{0})$ is defined by
@@
It was introduced in [51] (when $\Delta=0)$ [^3] In the notation of [51], $\mathcal{D}_{\gamma}(u)=F_{\gamma}(u),$ but here the notation $F_{\gamma}$ is reserved for the free energy functional.. Here we will consider the case of a log Fano curve and allow $u\in C^{1}(X)$ (thus, importantely, we do not assume $\omega_{u}\geq 0).$ In this case $\mathcal{E}(u)$ is given by the classical expression in Example 3.2. We will restrict $\mathcal{D}_{\gamma}(u)$ to the subspace $C^{1}(X)_{0}$ of all $u\in C^{1}(X)$ satisfying
@@
This is essentially shown in [10], but for completeness we recall the proof. The starting point is that when $n=1$, the sup in the definition 3.6 of $E(\mu)$ may be taken over all of $C^{0}(X),$ as pointed out in Example 3.2. Hence, $E(\mu)\leq\left\langle v,\mu\right\rangle-\mathcal{E}(v)$ for any $v\in C^{0}(X).$ Now take $\mu=e^{-\gamma u}/\int_{X}e^{-\gamma u}\mu_{0}$ and apply the previous inequality for $v=u.$ The inequality 4.6 then follows from the identity
@@
which is a direct consequence of the definition of $D(\mu|\mu_{0}).$ Thanks to the inequality 4.6 the lower bound on $\mathcal{D}_{\gamma}$ in the corollary thus follows directly from the lower bound on $\delta^{A}(X,\Delta_{w})^{\mathcal{G}}$ in Theorem 3.13. Next, to destabilize $\mathcal{D}_{\gamma}(u),$ first recall [10] that for any $u\in\mathcal{H}(X,\omega_{0})$
@@
In fact, this inequality holds also when $n>1$ by replacing $\omega_{u}$ with $\omega_{u}^{n}.$ Indeed, $F_{\gamma}(\frac{\omega_{u}}{V})-\mathcal{D}_{\gamma}(u)=D(\frac{\omega_{u}^{n}}{V}|\frac{\mu_{\phi_{0}+u}}{\int\mu_{\phi_{0}+u}})\geq 0,$ using that $D(\cdot|\cdot)\geq$ on $\mathcal{P}(X)^{2}.$ All that remains is thus to show that the curve $\mu_{t}$ in $\mathcal{P}(X)_{0}$ destabilizing $F_{-\gamma},$ constructed in the proof of Theorem 3.13, has the property that it can be expressed as $\mu_{t}=\omega_{u_{t}}/V$ for $u_{t}$ in $\in C^{1,1}(X)_{0}\cap\text{PSH}(X,\omega_{0}).$ To this end first consider the case when $w<1/2,$ i.e. $\min\{\frac{1}{1-w},2\}<2.$ Then $u_{t}:=\phi_{t}-\phi_{0},$ where $\phi_{t}$ is defined by 3.39, has the desired properties. Indeed, $\mu_{t}$ destabilizes $F_{-\gamma}$ on $\mathcal{P}(X)_{0}$ and, just as before, the $\mathbb{Z}_{2}$–invariance of $u_{t}$ (resulting from the $\mathbb{Z}_{2}$–invariance of $\phi_{t}$ and $\phi_{0})$ implies that $e^{-\gamma u_{t}}\mu_{0}$ has vanishing moment. Finally, when $w\geq 1/2$ the curve $u_{t}:=\phi_{t}-\phi_{0},$ where $\phi_{t}$ is the Legendre transform of $v_{t},$ defined by formula 3.37, has the desired properties. Again, this follows from the symmetry argument employed in the proof of Theorem 3.13, since $u_{t}$ is $\mathbb{Z}_{2}$–invariant. ∎
@@
is equivalent to the following Moser–Trudinger inequality with Aubin’s constraint [^4] Strictly speaking, Aubin’s inequalities [7, Cor 2] are obtained by replacing $\varphi$ with $\varphi$: